A researcher wishes to see if the average number of sick days a worker takes per year is greater than 2.8. A random sample of 30 workers at a large department store had a mean of 3.1. The standard deviation of the population is 0.8. Is there enough evidence to support the researchers claim at α = 0.01?
Answer #1

This hypothesis test is a right tailed test .

The null and alternative hypothesis is

H0 : \mu = 2.8

Ha : \mu > 2.8

Test statistic = z

= (\bar x - \mu ) / \sigma    / \sqrt n

= (3.1 - 2.8) / 0.8 / \sqrt 30

Test statistic = 2.054

P-value = 0.02

P-value > \alpha

Fail to reject the null hypothesis .

There is not enough evidence to support the researchers claim at α = 0.01 .